# ASTRID - An Automated and Scalable TRIaD for the Evaluation of RAG-based Clinical Question Answering Systems

### v1

@@ A.5 Illustrative Examples of ASTRID
A.6 Dataset Curation Process
To collect real-world patient questions, we used a autonomous telemedicine assistant capable of conducting phone conversations and answering patient questions regarding their recovery following cataract surgery. From these interactions, we gathered 102 unique questions from 120 patients from calls that took place as a standard of their care at two UK hospitals. All patients explicitly consented to the use of their anonymised data for research purposes.
To generate answers to these questions, we curated a knowledge source on cataract surgery with the help of two ophthalmic surgeons. We then employed three LLMs – Palm-2 (text-bison@002, [1]), Mistral-7B [17]) and Llama-3-8B [44] – as part of a RAG-based QA agent to generate responses to the 102 questions. This process resulted in a dataset of 306 question-answer-context triplets.
@@ A.6 Dataset Curation Process
We acknowledge that dataset size is limited as real-world clinical data is expensive to gather, and using simulated data may have invalidated some of the claims of connecting our triad with real-world safety performance of automated metrics. We plan to extend our evaluations with larger datasets in future work to reinforce our findings.
A.6.1 Balancing by Perceived Faithfulness
Two labellers assessed faithfulness for the 206 examples by showing them only the answer and the context. We asked them to use their own judgement to determine whether a given answer was faithful to the context. We refer to this measure of human judgment as Perceived Faithfulness (PF). The labellers discussed and resolved any disagreements to ensure consensus.
@@ A.6.1 Balancing by Perceived Faithfulness
A.6.2 Augmenting with out-of-scope data
For a holistic evaluation, we augmented this dataset with 45 out-of-scope questions selected by two clinicians from the open-source dataset HealthSearchQA [37]. We created 90 question-answer-context triplets using the same process mentioned earlier with only Palm-2 and Llama-3-8B, resulting in a comprehensive dataset of total 238 triplets.
A.6.3 Human labelling
Two labellers were asked to follow the older definition of Faithfulness and Conversational Faithfulness (CF) to generate labels for the (answer-context) pairs from the 238 examples. After the labellers resolved disagreements, we combined these ratings with the perceived faithfulness human ratings to create a comprehensive dataset of 238 question-answer-context triplets with human-rated faithfulness, CF, and PF. We name this dataset FaithfulnessQAC and make it available for research purposes.
A.6.4 Clinician Labelling
Clinician labellers were provided with the context of the task, as well as labels with definitions as per Figure 12.
Figure 12: Clinical Label Definitions for harm labelling.
A.6.5 Creating a dataset of unique questions
@@
A.6.6 Creating a dataset of unique questions
To support Claim (2), we need a dataset where each question is unique, as Context Relevance is measured between the question and the context. To achieve this, we sampled 87 out of 148 triplets where each question is unique, with the corresponding answers randomly selected from one of the three LLM responses. For the out-of-scope data, we randomly chose 45 triplets with responses from between the two models. The final dataset contains 132 question-answer-context triplets, along with the human ratings for CF, CR and RA from labellers. We name this dataset UniqueQAC and make it available for research purposes.
A.6.7 Creating a dataset for clinical harm, helpfulness and inappropriateness
@@
    Inappropriateness: Is there inappropriate or incorrect content?
@@
Figure 14: ClinicalQAC test split distribution across categories
A.7 Experiment Details
@@
A.7.1 System Setup
For all training and data analysis, we use Google Colaboratory’s unpaid version. For computing metrics, we ran the code on MacBook Pro M3. We host Palm-2, Gemini, Llama-3-8B, Llama-3-70B, Mistral-7B, Mistral-8x7B, and Claude Opus via Google’s Vertex AI Platform. We signed an agreement for Claude Opus via Vertex AI seeking permission to use it for research purposes.
@@ A.7.1 System Setup
We implement a random forest classifier using Scikit-learn. We perform grid on the parameters n_estimators, max_depth, min_samples_split, min_samples_leaf and bootstrap.
@@
We implement an SVM using Scikit-learn. We perform grid on the parameters C, gamma and kernel, except for Helpfulness.
@@
We implement a Gaussian Naive Bayes model using Scikit-learn.
@@
We implement a simple neural network using Pytorch. We use Cross Entropy loss and Adam optimiser.
1 class SimpleNN(nn.Module):
2 def __init__(self, input_size, hidden_size, output_size):
3 super(SimpleNN, self).__init__()
4 self.fc1 = nn.Linear(input_size, hidden_size)
5 self.relu = nn.ReLU()
6 self.fc2 = nn.Linear(hidden_size, hidden_size)
7 self.fc3 = nn.Linear(hidden_size, output_size)
8 self.softmax = nn.Softmax(dim=1)
9
10 def forward(self, x):
11 out = self.fc1(x)
12 out = self.relu(out)
13 out = self.fc2(out)
14 out = self.relu(out)
15 out = self.fc3(out)
16 out = self.softmax(out)
17 return out
18
19 input_size = X_train.shape[1]
20 hidden_size = 16
21 output_size = len(label_encoder.classes_)
22 model = SimpleNN(input_size, hidden_size, output_size)
A.7.2 Harmfulness
Table 4: Best Hyperparameters for Random Forest Classifier used for Harmfulness
Hyperparameter Value
bootstrap True
max_depth None
min_samples_leaf 1
min_samples_split 2
n_estimators 100
Table 5: Best Hyperparameters for SVM used for Harmfulness
Hyperparameter Value
C 10
gamma 0.1
kernel RBF
Figure 15: Training curves for Harmfulness
A.7.3 Helpfulness
Table 6: Best Hyperparameters for Random Forest Classifier used for Helpfulness
Hyperparameter Value
bootstrap True
max_depth None
min_samples_leaf 1
min_samples_split 2
n_estimators 200
Figure 16: Training curves for Helpfulness
A.7.4 Inappropriateness
Table 7: Best Hyperparameters for Random Forest Classifier used for Inappropriateness
Hyperparameter Value
bootstrap True
max_depth None
min_samples_leaf 1
min_samples_split 2
n_estimators 100
Table 8: Best Hyperparameters for SVM used for Inappropriateness
Hyperparameter Value
C 10
gamma 1
kernel RBF
Figure 17: Training curves for Inappropriateness
A.8 Inter-annotator agreements
@@
    Intent: Does it address the intent of the question?
@@
Table 9: Inter-annotater agreement on clinical axes

### v2

@@ Appendix D Illustrative Examples of ASTRID
Appendix E Dataset Curation Process
To collect real-world patient questions, we used an autonomous telemedicine assistant capable of conducting phone conversations and answering patient questions regarding their recovery following cataract surgery. From these interactions, we gathered 102 unique questions from 120 patients from calls that took place as a standard of their care at two UK hospitals. All patients explicitly consented to the use of their anonymised data for research purposes.
To generate answers to these questions, we curated a knowledge source on cataract surgery with the help of two ophthalmic surgeons. We then employed three LLMs – Palm-2 (text-bison@002, Anil et al. (2023)), Mistral-7B [Jiang et al. (2023)) and Llama-8B [Touvron et al. (2023)] – as part of a RAG-based QA agent to generate responses to the 102 questions. This process resulted in a dataset of 306 question-answer-context triplets.
@@ Appendix E Dataset Curation Process
E.1 Balancing by Perceived Faithfulness
We employed two independent labellers to assess "faithfulness" for the 206 examples by showing them only the answer and the context. We asked them to use their own judgment to determine whether a given answer was faithful to the context. We refer to this measure of human judgment as Perceived Faithfulness (PF). The labellers discussed and resolved any disagreements to ensure consensus.
@@ E.1 Balancing by Perceived Faithfulness
E.2 Augmenting with out-of-scope data
For a holistic evaluation, we augmented this dataset with 45 out-of-scope questions selected by two clinicians from the open-source dataset HealthSearchQA [Singhal et al. (2023a)]. We created 90 question-answer-context triplets using the same process mentioned earlier with only Palm-2 and Llama-8B, resulting in a comprehensive dataset of total 238 triplets.
E.3 Human labeling
We employed two independent labellers to follow the older definition of Faithfulness and Conversational Faithfulness (CF) to generate labels for the (answer-context) pairs from the 238 examples. After the labellers resolved disagreements, we combined these ratings with the perceived faithfulness human ratings to create a comprehensive dataset of 238 question-answer-context triplets with human-rated faithfulness, CF, and PF. We name this dataset FaithfulnessQAC and make it available for research purposes.
E.4 Creating a dataset of unique questions
@@
E.5 Creating a dataset of unique questions
To support Claim (2), we need a dataset where each question is unique, as Context Relevance is measured between the question and the context. To achieve this, we sampled 87 out of 148 triplets where each question is unique, with the corresponding answers randomly selected from one of the three LLM responses. For the out-of-scope data, we randomly chose 45 triplets with responses from between the two models. The final dataset contains 132 question-answer-context triplets, along with the human ratings for CF, CR and RA from independent labellers. We name this dataset UniqueQAC and make it available for research purposes.
E.6 Creating a dataset for clinical harm, helpfulness and inappropriateness
@@
    Appropriateness: Is there inappropriate or incorrect content?
@@
Appendix F Experimental details
@@
We implement a random forest classifier using Scikit-learn. We perform grid on the parameters
@@
We implement an SVM using Scikit-learn.
@@
We implement an Gaussian Naive Bayes using Scikit-learn.
@@
We implement a simple neural network using PyTorch.
Appendix G Inter-annotator agreements
@@
    Intent: Does it address the question intent?
@@
Table 4: Inter-annotator agreement on clinical axes
Appendix H Additional LLM Performance Analysis on ASTRID Metrics
To provide a more detailed view of model performance across the ASTRID metrics, Figures 17, 18, and 19 present Accuracy, F1-score, Precision, and Recall for each model across Conversational Faithfulness, Context Relevance, and Refusal Accuracy.
Figure 17: Model-wise performance on Conversational Faithfulness across Accuracy, F1-score, Precision, and Recall.
Figure 18: Model-wise performance on Context Relevance across Accuracy, F1-score, Precision, and Recall.
Figure 19: Model-wise performance on Refusal Accuracy across Accuracy, F1-score, Precision, and Recall.
Model Provider Date Temperature Top p Token Limit
llama3-3-70b-instruct AWS 12/02/2025 0.1 0.9 200
llama3-8b-instruct AWS 12/02/2025 0.1 0.9 200
mistral-7b-instruct AWS 12/02/2025 0.1 0.9 200
mistral-large AWS 12/02/2025 0.1 0.9 200
gemini-2.0-flash Google 12/02/2025 0.1 0.9 200
claude-3.5-sonnet Google 12/02/2025 0.1 0.9 200
gpt-4o Azure 12/02/2025 0.1 0.9 200
gpt-o3-mini Azure 12/02/2025 N/A N/A 1000
Table 5: Inference details for all evaluated LLMs.